# Understanding In-context Learning of Addition via Activation Subspaces

### v2

@@ Abstract
To perform in-context learning, language models must extract signals from individual few-shot examples, aggregate these into a learned prediction rule, and then apply this rule to new examples. How is this implemented in the forward pass of modern transformer models? To study this, we consider a structured family of few-shot learning tasks for which the true prediction rule is to add an integer $k$ to the input. We find that Llama-3-8B attains high accuracy on this task for a range of $k$, and localize its few-shot ability to just three attention heads via a novel optimization approach. We further show the extracted signals lie in a six-dimensional subspace, where four of the dimensions track the unit digit and the other two dimensions track overall magnitude. We finally examine how these heads extract information from individual few-shot examples, identifying a self-correction mechanism in which mistakes from earlier examples are suppressed by later examples. Our results demonstrate how tracking low-dimensional subspaces across a forward pass can provide insight into fine-grained computational structures.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable in-context learning (ICL) capabilities, where they can adapt to various new tasks through few-shot demonstrations in the prompt. Motivated by this, researchers have constructed detailed models of in-context learning for small synthetic language models (Garg et al., 2022; Akyürek et al., 2023; von Oswald et al., 2023) as well as coarser-grained treatments of large pretrained models (Olsson et al., 2022; Hendel et al., 2023; Todd et al., 2024). However, little is known about the fine-grained computational structure of in-context learning for large models.
To address this, we perform a detailed analysis of ICL for a structured family of tasks: learning to add a constant number (“add-k”). For these tasks, models are given an input sequence “$x_{1}\rightarrow y_{1}\#...\#x_{n}\rightarrow y_{n}\#x_{q}\rightarrow$” where $y_{i}=x_{i}+k$ for all $i$. This task is mathematically well-defined and tractable for an in-depth mechanistic analysis, while exhibiting the core challenge of extracting and applying a pattern from ICL examples. Moreover, modern language models are empirically capable of solving it (they correctly predict $y_{q}=x_{q}+k$).
@@
Figure 1: In this paper, we study the activation space of attention heads in Llama-3-8B during in-context learning (ICL) for addition tasks. Our main findings: (1) out of $1024$ attention heads in Llama-3-8B, only three are important for these tasks (§4); (2) each head encodes the task information in a six-dimensional subspace (§4.1); (3) a six-dimensional subspace further decomposes into a four-dimensional subspace capturing the unit digit of the task information and a two-dimensional subspace capturing its ten digit (§4.3); and (4) task information is mostly extracted from a subset of tokens (§5). The figure illustrates these findings using an example add-k prompt.
To understand the mechanism behind in-context add-k, we study it in Llama-3-8B. We first discover that only a small number of model components are responsible for add-k: out of 1024 attention heads, only three do most of the work (§3). We find the task information (i.e., the number $k$) for each head is encoded in a six-dimensional subspace; zooming into each head, we further decompose it into a four-dimensional space tracking the unit digit using sine and cosine functions and a two-dimensional one tracking the overall magnitude (§4). We also examine how this head aggregates information from individual in-context examples (§5).
To discover the attention heads that are responsible for add-k, we introduce a novel approach inspired by previous work on function vectors (Todd et al., 2024), which are vectors that imitate ICL when they are patched into the residual stream (§2.2). We formulate an optimization method that identifies $33$ important heads whose outputs yield good ICL performance (§3.1). Through further ablation, we find that all but 3 of these 33 heads can be replaced with their mean value without significantly decreasing the performance (§3.2).
Since these three heads are responsible for computing add-k, they must extract signal about $k$ from individual examples and then represent $k$ in the residual stream. To understand how $k$ is represented, zooming into each head, we first find that the effect of each head factors through only a six-dimensional subspace (out of $128$ in total) via principal component analysis (PCA) (§4.1). Furthermore, we find that it factors into a four-dimensional subspace tracking the unit digit through trigonometric functions, and a two-dimensional subspace tracking the coarse-grained magnitude (§4.2-4.3). This offers insight on how LLMs internally represent addition, aligning with recent findings showing that (non-ICL) addition is encoded via trigonometric (Zhou et al., 2024) or helical functions (Kantamneni & Tegmark, 2025).
Finally, we examine how this head aggregates information from individual in-context examples. We find that the information is mostly extracted from the $y_{i}$ token positions, and that this extracted signal encodes a vector representation of $y_{i}-x_{i}$5.2). Moreover, we identify a self-correction mechanism: if the signal extracted from one token has an error, later tokens are likely to write in the opposite direction of the error (§5.3). This means that ICL goes beyond simple averaging of inputs and has stateful dynamics.
Our findings suggest that ICL tasks can be implemented through specialized activation subspaces from a small number of attention heads that extract, represent and aggregate task-related information in interpretable ways. We introduce a novel optimization method to identify important attention heads and demonstrate how a focused analysis of a simple ICL task can reveal delicate latent space structures and sophisticated computational strategies in a large language model.
@@ 2.1 Model and Task
Model. We study Llama3-8B-Instruct, which has 32 layers and 32 attention heads per layer, and a residual dimension of 4096. We denote each head as a tuple (layer index, head index), where both indices range from 0 to 31.
@@
For a constant $k$, the add-k task is to add $k$ to a given input integer $x$ to predict $y=x+k$. In an $n$-shot ICL prompt, we provide $n$ examples of the format “$x_{i}\rightarrow y_{i}$” (where each $y_{i}=x_{i}+k$). The examples are concatenated in the prompt using the separator “##\##”, followed by a query “$x_{q}\rightarrow$”; see Figure 1 for an example of a five-shot ICL prompt. A key advantage of this family of tasks is that ICL prompts for different tasks only differ in $k$ (i.e. $y_{i}-x_{i}$ in the prompts) but not the input domain, in contrast to most previous studies (Todd et al., 2024; Hendel et al., 2023).
We construct data for the task by varying $x\in[1,100]$ and $k\in[1,30]$ (thus $y\in[2,130]$) since Llama3-8B is empirically capable of solving the addition task in this range. We consider the following three types of prompts:
@@
    five-shot ICL prompt, where all 5 demonstration examples satisfy $y_{i}=x_{i}+k$ for a fixed $k\in[1,30]$. We also call this add-k ICL.
@@
    mixed-$k$ ICL prompt, where the five demonstration examples are $y_{i}=x_{i}+k_{i}$ for five possibly different $k_{i}$ values.
@@
    zero-shot prompt, where there are no demonstration examples: the prompt is “$x_{q}\rightarrow$” for some $x_{q}$.
@@ 2.2 Activation Patching and Function Vectors
Next, we briefly introduce activation patching, a common interpretability technique that will be used throughout the paper, and the function vector, a construction that we will use to identify important heads for our ICL tasks.
@@
Specifically, if $z_{l}$ is the original value of the residual stream at layer $l$ at the final token position $\to$”, we patch in the replacement $z_{l}+v$, where $v$ is constructed from activations on a different input prompt. We choose the layer $l$ at one third of the network’s depth and construct $v$ from “function vector” heads, following Todd et al. (2024), as described next.
@@
Function vectors $v_{k}$ are vectors constructed from the outputs of selected attention heads, designed so that adding $v_{k}$ to the residual stream of a zero-shot prompt approximates the effect of the five-shot add-k task. For example, $v_{k}$ might be the average output of one or more attention heads across a set of five-shot add-$k$ examples. We define the intervention accuracy of $v_{k}$ as the average accuracy obtained on zero-shot prompts when $v_{k}$ is added to the residual stream across all tasks $k$. This metric captures how effectively $v_{k}$ encodes task-specific information about $k$. For comparison, we define clean accuracy as the accuracy on five-shot prompts without any intervention, also averaged across all tasks $k$.
In more detail, consider an attention head $h$: let $h(p)$ denote the output of head $h$ on prompt $p$ at the last token position, and define $h_{k}$ as the average of $h(p)$ across all five-shot add-k prompts. [^1] Similar to Todd et al. (2024), we approximate this average using $100$ randomly generated five-shot add-k prompts, where each prompt contains random demonstration examples $x_{i}$ and the final query input $x_{q}$ is chosen exactly once from each integer in $[1,100]$. Todd et al. (2024) identified a subset $\mathcal{H}$ of attention heads (for a different set of tasks) such that the vector $v_{k}=\sum_{h\in\mathcal{H}}h_{k}$ has high intervention accuracy—that is, adding it to the residual stream of zero-shot prompts effectively recovers few-shot task performance.
Building on this framework, we consider multiple ways to construct $v_{k}$: (1) the task-specific mean $h_{k}$ over the add-k task (as described above); (2) the overall mean $\bar{h}$ (i.e., average across all $k$), which removes task-specific information about $k$; or (3) the specific value $h(p)$ on an individual prompt $p$. Throughout the paper, we call $\bar{h}$ the mean-ablation of head $h$ and call $h_{k}$ the head vector of $h$. Beyond the three choices above, we sometimes project a head’s output onto a lower-dimensional subspace or scale it by a coefficient. Unlike Todd et al. (2024), we recover a different set of heads through solving a novel global optimization problem (§3.1) and systematic ablation studies (§3.2), which achieves higher performance (Figure 3).
Figure 2: The chain of localization in §3 and §4. We first identify $33$ significant attention heads (out of 1024) via a global optimization method (§3.1), then narrow down to 3 main heads while mean-ablating the remaining 30 heads (§3.2). We next study the structure of the representations of each main head by localizing it to a six-dimensional subspace (§4.1), and decompose it into a four-dimensional subspace encoding the unit digit and a two-dimensional subspace encoding the tens digit (§4.3).
@@ 3 Identifying Three Aggregator Heads
To understand the mechanism behind the add-k ICL task, we first need to find out what model components are responsible for performing it. In this section, we find that three heads do most of the work for add-k. We find these heads by first solving an optimization problem to identify $33$ significant heads out of 10241024102410243.1), and then further narrowing down to three main heads via systematic ablations (§3.2). This process is illustrated in the first row of Figure 2.
@@ 3.1 Identifying Significant Heads via Sparse Optimization
Let $\ell(x_{q},y_{q};v)$ be the cross-entropy loss between $y_{q}$ and the model output when intervening the vector $v$ onto the input “$x_{q}\to$”(i.e. replacing $z_{l}$ with $z_{l}+v$ in the forward pass on “$x_{q}\to$”, where $z_{l}$ is the layer-$l$ residual stream at the last token). We optimize $c$ with respect to the loss
$\mathcal{L}(c)=\mathbb{E}_{k\in[30]}\mathbb{E}_{x_{q}\in[100]}[\ell(x_{q},x_{q}+k;v_{k}(c))]+\lambda\|c\|_{1},$ (1)
@@
(a) Coefficients of heads.
(b) Average indirect effects of heads (Todd et al., 2024).
Figure 3: Comparison of significant heads identified by our optimized coefficients (left) and by average indirect effects (AIE) from the previous method (Todd et al., 2024) (right). Colors indicate the magnitude of each head’s importance (coefficients or AIE). The top $33$ heads identified by both methods are highlighted with frames ($13$ heads common across both methods in red and other $20$ heads in blue). Our identified heads yield an intervention accuracy of $0.85$, compared to the previous method’s accuracy of $0.31$. Both methods select heads from similar layers, but our optimization approach is significantly more effective.
Training details. We randomly select $25$ add-k tasks from our total $30$ tasks for training and in-distribution testing, and use the remaining five tasks only for out-of-distribution testing. For each task add-k, we generate $100$ zero-shot prompts “$x_{q}\rightarrow$”, where $x_{q}$ ranges over all integers from $[1,100]$ and the target output is $x_{q}+k$, yielding one data point for each (prompt, task) pair. We randomly split the data points of the $25$ tasks into training, validation, and test sets in proportions of $0.7$, $0.15$, and $0.15$, respectively. We use AdamW with learning rate $0.01$ and batch size $128$. We set the regularization rate $\lambda$ as $0.05$, which promotes sparsity while incurring little loss in accuracy. During training, we clip the coefficients $c$ back to $[0,1]$ if they go out of range after each gradient step.
Results. We get coefficients that achieves high intervention accuracy. In particular, at the final epoch, the average accuracies of intervening on test data for the $25$ tasks and the five tasks ($0.83$ and $0.87$) are both close to the average accuracies of doing no intervention ($0.89$ and $0.92$). To identify the important heads for the tasks, we plot the values of the coefficients in the final epoch for each layer and head index (Figure 3(a)). We find $33$ heads have coefficients $c_{h}$ greater than $0.2$, most ($21$) of which are one. In contrast, the other ($991$) coefficients are all smaller than $0.01$, most ($889$) of which are zero. We call the $33$ heads significant heads and denote the set of them as $\mathcal{H}_{\text{sig}}$.
Comparison to previous approach. We compare our optimization approach with the previous method from Todd et al. (2024) for identifying important heads, which selects heads based on average indirect effect (AIE). To perform a fair comparison, we construct our function vector by summing the outputs of our selected heads directly without weighting by coefficients, matching their methodology. Using this construction, we achieve an intervention accuracy of $0.85$, close to the clean accuracy of $0.87$, indicating that our $33$ heads captures most of the necessary information for the add-k task. In contrast, selecting the top 33 heads according to AIE yields a much lower intervention accuracy of $0.31$. [^2] Todd et al. select the top 10 heads in their work, but this yields an even lower accuracy of $0.05$ in our setting. In Figure 3, we visualize the coefficients and AIE values of heads from both methods and highlight the $13$ heads that are common to both approaches.

### v3

@@ Abstract
To perform few-shot learning, language models extract signals from a few input-label pairs, aggregate these into a learned prediction rule, and apply this rule to new inputs. How is this implemented in the forward pass of modern transformer models? To explore this question, we study a structured family of few-shot learning tasks for which the true prediction rule is to add an integer $k$ to the input. We introduce a novel optimization method that localizes the model’s few-shot ability to only a few attention heads. We then perform an in-depth analysis of individual heads, via dimensionality reduction and decomposition. As an example, on Llama-3-8B-instruct, we reduce its mechanism on our tasks to just three attention heads with six-dimensional subspaces, where four dimensions track the unit digit with trigonometric functions at periods $2$, 55, and $10$, and two dimensions track magnitude with low-frequency components. To deepen our understanding of the mechanism, we also derive a mathematical identity relating “aggregation” and “extraction” subspaces for attention heads, allowing us to track the flow of information from individual examples to a final aggregated concept. Using this, we identify a self-correction mechanism where mistakes learned from earlier demonstrations are suppressed by later demonstrations. Our results demonstrate how tracking low-dimensional subspaces of localized heads across a forward pass can provide insight into fine-grained computational structures in language models.
@@ 1 Introduction
Large language models (LLMs) exhibit in-context learning (ICL) abilities; for instance, they can few-shot learn new tasks from a small number of demonstrations in the prompt. To understand this ability, past works have constructed detailed models of ICL for small synthetic language models (Garg et al., 2022; Akyürek et al., 2023; von Oswald et al., 2023) as well as coarser-grained analyses of large pretrained models (Olsson et al., 2022; Hendel et al., 2023; Todd et al., 2024). However, little is known about the fine-grained computational structure of ICL for large models.
ICL extracts task information from demonstrations and applies the aggregated information to input queries. Previous work (Todd et al., 2024) constructed a vector (i.e., the “function” vector) from demonstrations that encode the task information—for instance, adding it to the residual stream on zero-shot inputs recovers ICL behavior. However, two questions remain elusive: (1) How precisely do function vectors encode task information? (2) How do models aggregate information from ICL examples to form such function vectors?
To address these questions, we perform a case study for few-shot learning of arithmetic (i.e., learning to add a constant $k$ to the input). This family of tasks has the advantage of providing a large number of tasks (different integers $k$) that all share the same input domain, which is important to rule out domain-based shortcuts when analyzing ICL mechanisms (§2.1). Arithmetic also has the advantage of being well-studied in other (non-in-context) settings (Zhou et al., 2024; Kantamneni & Tegmark, 2025), allowing us to situate our results with other known mechanisms. Finally, most LLMs perform this task reliably (e.g. $87\%$ accuracy for Llama-3 and $90\%$ for Qwen-2.5).
To study this setting, we first introduce a novel optimization method for localizing few-shot ability to a small number of attention heads. Our approach is inspired by previous work on function vectors (Todd et al., 2024), which mimic ICL when they are patched into the residual stream (§2.2). While Todd et al. (2024) selected heads based on their individual effect on ICL performance, we search for coefficients within a continuous range $[0,1]$ that produce a sparse, weighted combination of heads’ outputs, maximizing ICL performance (§3.1). Our method successfully finds a small number of heads that recover most of the ICL performance (§3.2); for example, Llama-3-8B-instruct achieves 79% accuracy using function vectors from just three heads (90% of the original ICL accuracy).
We next perform a detailed analysis of these heads, through dimensionality reduction and decomposition. We find that the task information from each head is encoded in a low-dimensional subspace typically consisting of trigonometric functions. For example, in each of the three important heads in Llama-3-8B-instruct, we get a 6-dimensional subspace, where 4 dimensions are periodic functions with periods 2, 5, and 10 (tracking the units digit) and 2 dimensions are low-frequency functions (tracking the tens digit) (§4). This interestingly aligns with recent findings in non-in-context settings where addition is also encoded by trigonometric (Zhou et al., 2024) or helical functions (Kantamneni & Tegmark, 2025), suggesting a deeper relation between non-ICL and ICL machinery.
@@
Figure 1: Key findings of our methods in the specific case of Llama-3-8B-instruct (illustrated using an example add-k prompt): (1) out of 1024 attention heads, only three are important for add-k ICL (§4); (2) each head encodes the task information $k$ in a six-dimensional subspace (§4.1); (3) the six-dimensional subspace further factors into four dimensions capturing the unit digit of $k$ (encoding periodic functions at periods 2,5, and 10) and two dimensions capturing its tens digit (encoding higher frequency functions) (§4.3); and (4) task information flows from output tokens to final token via a self-correction mechanism (§5).
Finally, we further study the flow of information across tokens, by deriving a general mathematical relation between “extractor” and “aggregator” subspaces, building on the second-order logit lens (Gandelsman et al., 2025). This lets us study the task-related information that heads extract from each token. For example, in Llama-3-8B-instruct we find that task information is mostly extracted from label tokens (§5.2). Moreover, we find a self-correction mechanism: if the signal extracted from one token has an error, signals from later tokens often write in the opposite direction of the error (§5.3). This suggests that ICL goes beyond simple averaging of inputs and has stateful dynamics.
In summary, we found that task-specific information in ICL can be localized to a small number of attention heads and low-dimensional subspaces with structured patterns. We also found that models employ a self-correction mechanism when the task information flows from the “extractor” subspaces to the “aggregator” subspaces. Our findings show how even in large models, ICL mechanisms can be localized to specialized activation subspaces from a small number of heads that extract, represent, and aggregate information in interpretable ways.
Our key contributions are thus as follows: (1) we introduce a novel optimization method to identify relevant attention heads for ICL; (2) we derive a precise mathematical relation between signals from earlier tokens and the output token; and (3) we perform a focused analysis of ICL addition and are able to reverse-engineer rich latent structures and sophisticated computational strategies in LLMs.
@@ 2.1 Model and Task
Model. We focus on Llama-3-8B-instruct in the main body, which has 32 layers and 32 attention heads per layer and a residual dimension of 4096. We denote each head as a tuple (layer index, head index), where both indices range from 0 to 31. To make our analysis broader in model size, training type, and model family, we also experiment on Llama-3.2-3B-instruct, Llama-3.2-3B, and Qwen-2.5-7B and report the results in appendix. The results are consistent across Llama family and directionally similar for the Qwen model.
@@
We study a structured family of ICL tasks that we call add-$k$. For a constant $k$, the add-$k$ task is to add $k$ to a given input integer $x$ to predict $y=x+k$. In an $n$-shot ICL prompt, the model is given $n$ demonstrations of the form $x_{i}\rightarrow y_{i}$” with $y_{i}=x_{i}+k$, concatenated using the separator “#\#”, followed by a query $x_{q}\rightarrow$” (see Figure 1 for an example). A key advantage of this family of tasks is that ICL prompts for different tasks only differ in $k$ (i.e. $y_{i}-x_{i}$) but not in the input domain, enabling us to isolate task information from input content so as to dissect the ICL mechanism at a finer granularity.
[…]
We construct data for the task by varying $x\in[1,100]$ and $k\in[1,30]$ (thus $y\in[2,130]$) since Llama-3 models are empirically capable of solving the addition task in this range. We consider the following three types of prompts:
@@
    five-shot ICL prompt, where all five demonstrations satisfy $y_{i}=x_{i}+k$ for a fixed $k\in[1,30]$. We also call this add-k ICL.
@@
    mixed-$k$ ICL prompt, where the demonstrations are $y_{i}=x_{i}+k_{i}$ for possibly different $k_{i}$ values.
@@
    zero-shot prompt, where there are no demonstrations: the prompt is $x_{q}\rightarrow$” for some $x_{q}$.
We mostly study five-shot ICL prompts as examples of our ICL tasks, and use them to generate function vectors (§2.2). We also examine the information extracted from demonstrations in mixed-$k$ ICL prompts, which is a varied version of five-shot ICL prompts with mistaken demonstrations, in §5.2. We use zero-shot prompts to evaluate the effectiveness of heads and function vectors (§2.2).
@@ 2.2 Activation Patching and Function Vectors
Next, we briefly review activation patching, a common interpretability technique that will be used throughout the paper, and the function vector, a construction that we will use to identify important heads for our ICL tasks.
@@
Specifically, if $z_{l}$ is the original value of the residual stream at layer $l$ at the final token position $\to$”, we patch in the replacement $z_{l}+v$, where $v$ is constructed from activations on a different input prompt. We choose the layer $l$ at one-third of the network’s depth and construct $v$ from “function vector” heads, following Todd et al. (2024), as described next.
@@
Function vectors $v_{k}$ are vectors constructed from the outputs of selected attention heads, designed so that adding $v_{k}$ to the residual stream of a zero-shot prompt approximates the effect of the five-shot add-k task. For example, $v_{k}$ might be the average output of one or more attention heads across a set of five-shot add-$k$ examples. We define the intervention accuracy of $v_{k}$ as the average accuracy obtained on zero-shot prompts when $v_{k}$ is added to the residual stream across all tasks $k$. Due to the independency of $k$ from input queries, this metric captures how effectively $v_{k}$ encodes task-specific information about $k$. For comparison, we define clean accuracy as the accuracy on five-shot prompts without any intervention, also averaged across all tasks $k$.
In more detail, consider an attention head $h$: let $h(p)$ denote the output of head $h$ on prompt $p$ at the last token position, and define $h_{k}$ as the average of $h(p)$ across all five-shot add-k prompts. [^1] Similar to Todd et al. (2024), we approximate this average using $100$ randomly generated five-shot add-k prompts, where each prompt contains random demonstration inputs $x_{i}$ and the final query input $x_{q}$ is chosen exactly once from each integer in $[1,100]$. Todd et al. (2024) identified a subset $\mathcal{H}$ of attention heads (for a different set of tasks) such that the vector $v_{k}=\sum_{h\in\mathcal{H}}h_{k}$ has high intervention accuracy—that is, adding it to the residual stream of zero-shot prompts effectively recovers few-shot task performance.
Building on this framework, we consider multiple ways to construct $v_{k}$: (1) the task-specific mean $h_{k}$ over the add-k task (as described above); (2) the overall mean $\bar{h}$ (i.e., average across all $k$), which removes task-specific information about $k$; or (3) the specific value $h(p)$ on an individual prompt $p$. Throughout the paper, we call $\bar{h}$ the mean-ablation of head $h$ and call $h_{k}$ the head vector of $h$ with respect to $k$. Beyond the three choices above, we sometimes project a head’s output onto a lower-dimensional subspace or scale it by a coefficient. Unlike Todd et al. (2024), we recover a different set of heads through solving a novel global optimization problem (§3.1) and systematic ablation studies (§3.2), which achieves higher performance (Figure 3).
Figure 2: The chain of localization in §3 and §4. We first identify $33$ significant attention heads (out of 1024) via a global optimization method (§3.1), then narrow down to 3 main heads while mean-ablating the remaining 30 (§3.2). We next study the structure of the representation of each main head by localizing it to a six-dimensional subspace (§4.1), and decompose it into a four-dimensional subspace encoding the unit digit and a two-dimensional subspace encoding the tens digit (§4.3).
@@ 3 Identifying Three Aggregator Heads
To understand the mechanism behind the add-k ICL task, we first need to find out what model components are responsible for performing it. In this section, we find that three heads do most of the work for add-k. We find these heads by first solving an optimization problem to identify $33$ significant heads out of 102410243.1), and then further narrowing down to three main heads via systematic ablations (§3.2). This process is illustrated in the first row of Figure 2.
@@ 3.1 Identifying Significant Heads via Sparse Optimization
Let $\ell(x_{q},y_{q};v)$ be the cross-entropy loss between $y_{q}$ and the model output when intervening the vector $v$ onto the input $x_{q}\to$”(i.e. replacing $z_{l}$ with $z_{l}+v$ in the forward pass on $x_{q}\to$”, where $z_{l}$ is the layer-$l$ residual stream at the last token). We optimize $c$ with respect to the loss
$\mathcal{L}(c)=\mathbb{E}_{k,x_{q}}[\ell(x_{q},x_{q}+k;v_{k}(c))]+\lambda\|c\|_{1},$ (1)
@@
Training details. We randomly select $25$ add-k tasks of the total $30$ tasks for training and in-distribution testing, and use the remaining five tasks only for out-of-distribution testing. For each task add-k, we generate $100$ zero-shot prompts $x_{q}\rightarrow$”, where $x_{q}$ ranges over all integers from $[1,100]$ and the target output is $x_{q}+k$, yielding one data point for each (prompt, task) pair. We randomly split the data points of the $25$ tasks into training, validation, and test sets in proportions $0.7$, $0.15$, and $0.15$, respectively. We use AdamW with learning rate $0.01$ and batch size $128$. We set the regularization rate $\lambda$ as $0.05$, which promotes sparsity while incurring little loss in accuracy. During training, we clip the coefficients $c$ back to $[0,1]$ if they go out of range after each gradient step.
Results. We get coefficients that achieve high intervention accuracy. In particular, on the $25$ in-distribution tasks and five out-of-distribution tasks, intervention accuracies at the final epoch are $0.83$ and $0.87$, close to the clean accuracies of $0.89$ and $0.92$, respectively. To identify the important heads for the tasks, we plot the values of the coefficients in the final epoch for each layer and head index (Figure 3(a)). We find $33$ heads have coefficients $c_{h}$ greater than $0.2$, most ($21$) of which are one. In contrast, the other ($991$) coefficients are all smaller than $0.01$, most ($889$) of which are zero. We call the $33$ heads significant heads and denote the set of them as $\mathcal{H}_{\text{sig}}$.
Comparison to previous approach. We compare our optimization approach with the previous method from Todd et al. (2024) for identifying important heads, which selects heads based on average indirect effect (AIE). To perform a fair comparison, we construct our function vector by summing the outputs of our selected heads directly without weighting by coefficients, matching their methodology. Using this construction, we achieve an intervention accuracy of $0.85$, close to the clean accuracy of $0.87$, indicating that our $33$ heads captures most of the necessary information for the add-k task. In contrast, selecting the top 33 heads according to AIE yields a much lower intervention accuracy of 0.310.31.222Todd et al. select the top 10 heads in their work, but this yields an even lower accuracy of $0.05$ in our setting. We visualize the coefficients and AIE values of heads from both methods in Figure 3.